Renee is twice as old as Dana. Four years ago, Renee was three times as old as Dana. How old are Renee and Dana now?

Answer by ewatrrr(24785)

 
Hi,
Let x represent Dana's age ⇒ Renee is 2x
2x - 4 = 3(x-4)
2x - 4 = 3x-12
8 = x, Dana's age. Renee is 16

CHECKING our answer*** 4 yrs ago
12 = 3*4
Wish You the Best in your Studies.